OpenStudy (anonymous):

What is the length of the side BC of triangle ABC? A. 5 cm B. 8 cm C. 26 cm D. 32 cm

OpenStudy (albert0898):

Please provide measurements or a picture.

OpenStudy (anonymous):

Sorry, OP is making my computer super slow!! :(

OpenStudy (anonymous):

I keep getting different answers!

OpenStudy (albert0898):

I'll do it and come back to you with the answer and how to do it. You will need to use the Pythagorean Theorem I believe. Ok the Theorem is... \[a^2 + b^2 = c^2\] 3x + 2 * 3x + 2 = \[9x^2 + 4\] 4x - 6 * 4x - 6 = \[16x^2 + -36\] \[9x^2 + 16x^2 = 25x^2\] 4 + -36 = -32 Final Answer = \[\sqrt{25x^2 - 32}\]

OpenStudy (anonymous):

24.35?

OpenStudy (anonymous):

or 5.65?

OpenStudy (albert0898):

That's the simplest form there is. If you enter \[\sqrt{25x^2 - 32}\] into a calculator... you'll get a domain error.

OpenStudy (albert0898):

I never did this before, you must understand that I'm in the 8th grade and haven't encountered such problems, but it seems as if you must you the Pythagorus.

OpenStudy (anonymous):

Thanks for trying! :/

OpenStudy (albert0898):

You're Welcome!